# An example of the Jantzen filtration of a D-module

### v1

@@ 3.4.3. Step 3: !!!-pushforward
As in 3.4.2, this $\widetilde{\mathcal{U}}$-module has an $n$-step filtration by powers of $s$, and decomposes into a direct sum of $\widetilde{\mathcal{U}}$-submodules, each spanned by the set of monomials $\{x_{1}^{k}x_{2}^{\ell}s^{m}\}$ and $\{x_{1}^{a}\partial_{2}^{b}s^{m}\}$ such that $k+\ell=a-b$ is a fixed integer. For $k+\ell=a-b=\lambda$, this submodule is isomorphic to a deformed Verma module of highest weight $\lambda$. We illustrate the module corresponding to $\lambda=0$ in Figure 4
@@ 3.5. The monodromy filtration and the geometric Jantzen filtration
We begin by recalling monodromy filtrations in abelian categories, following [Del80, §1.6]. Given an object $A$ in an abelian category $\mathcal{A}$ and a nilpotent endomorphism $s:A\rightarrow A$, it follows from the Jacobson-Morosov theorem [Del80, Proposition 1.6.1] that there exists a unique increasing exhaustive filtration $\mu^{\bullet}$ on $A$ such that $s\mu^{n}\subset\mu^{n-2}$, and for $k\in\mathbb{N}$, $s^{k}$ induces an isomorphism $\mathrm{gr}_{\mu}^{k}A\simeq\mathrm{gr}_{\mu}^{-k}A.$ This unique filtration is called the monodromy filtration of $A$. The monodromy filtration $\mu^{\bullet}$ induces filtrations $J_{!}^{\bullet}$ on $\ker s$ and $J_{+}^{\bullet}$ on $\operatorname{coker}s$. These sub and quotient filtrations can be described explicitly in terms of powers of $s$. Namely, if we set
@@
to be the image filtration of $A$, then it is shown in [Del80, §1.6] that
@@
(Deformation direction) There are other choices of regular functions on $\widetilde{X}$ which we could have used in the construction of these filtrations. In particular, if $\lambda\in\mathfrak{h}^{*}$ is dominant and integral such that $\lambda(h)=n$ for $n\in\mathbb{Z}_{\geq 0}$, then the function $f_{\lambda}:(x_{1},x_{2})\mapsto x_{2}^{n}$ can be used to define an intermediate extension functor $\Xi_{f_{\lambda}}$ and corresponding Jantzen filtrations. For $\mathfrak{g}=\mathfrak{sl}_{2}(\mathbb{C})$, all choices lead to the same filtration. However, for other Lie algebras, Williamson has shown that this is not the case [Wil16].
@@ 3.6.1. The algebraic Jantzen filtration
Denote by
(3.100) $I(\lambda):=\operatorname{Hom}_{\mathcal{U}(\mathfrak{b})}(\mathcal{U}(\mathfrak{g}),\mathbb{C}_{\lambda})$
the corresponding dual Verma module.
@@
trivially to $\mathfrak{b}$. Here $h\in\mathfrak{h}$, $a\in A$, and $\varphi$ is as in (3.101).
The deformed dual Verma module corresponding to $\lambda$ is
(3.104) $I_{A}(\lambda):=\operatorname{Hom}_{\mathcal{U}(\bar{\mathfrak{b}})}(\mathcal{U}(\mathfrak{g}),A_{\lambda}).$
There is a canonical isomorphism [Soe08, Prop 2.12]
@@
Under this isomorphism, $1\in A$ distinguishes a canonical $(\mathfrak{g},A)$-bimodule homomorphism [^11] This canonical map can be explicitly constructed using the unique nondegenerate $A$-valued contravariant form on $M_{A}(\lambda)$. This is how the filtration was originally defined in the work of Jantzen [Jan79].
(3.106) $\psi_{\lambda}:M_{A}(\lambda)\rightarrow I_{A}(\lambda).$
@@
Pulling back the filtration above along the canonical map $\psi_{\lambda}$ (3.106) gives a $(\mathfrak{g},A)$-bimodule filtration of $M_{A}(\lambda)$.
@@
$M_{A}(\lambda)^{i}:=\left\{m\in M_{A}(\lambda)\mid\psi_{\lambda}(m)\in s^{i}I_{A}(\lambda)\right\},$
where $\psi_{\lambda}$ is the canonical map (3.106). By applying the reduction map (3.108) to the filtration layers, we obtain a filtration $M(\lambda)^{\bullet}$ of $M(\lambda)$.
@@ 3.6.2. Relationship between algebraic and geometric Jantzen filtrations
As illustrated in Figures 3 and 4 , the global sections of $j_{!}f^{s}\mathcal{O}_{U}^{(n)}$ and $j_{+}f^{s}\mathcal{O}_{U}^{(n)}$ decompose into direct sums of deformed dual Verma and Verma modules, respectively. [^12] To be more precise, the submodules of $\Gamma(\widetilde{X},j_{!}f^{s}\mathcal{O}_{U}^{(n)})$ and $\Gamma(\widetilde{X},j_{+}f^{s}\mathcal{O}_{U}^{(n)})$ corresponding to an integer $\lambda$ are truncated versions of $M_{A}(\lambda)$ (3.104) and $I_{A}(\lambda)$ (3.104) obtained by taking a quotient so that $s^{n}=0$.. The global sections of $\mathrm{can}$ are the direct sum of $\psi_{\lambda}$ (3.106) for all integral $\lambda$. There are two natural filtrations of $j_{!}\mathcal{O}_{U}$ which we have described using this set-up.
@@
This is the geometric Jantzen filtration. It can be realized explicitly in terms of the image of powers of $s$ using 3.95. On global sections, this gives
@@
The map $\mathrm{can}$ is described on basis elements in (3.86). Computing these actions for $\lambda=0$, we see in Figure 8 that $\mathrm{can}$ fixes the right-most column and sends any other monomial on the left to a linear combination of monomials directly above the corresponding monomial on the right. The image of $s_{1}(n)=s\circ\mathrm{can}$ (3.89) is highlighted in grey. The quotient by this image is the maximal extension, which is outlined in the red box. The quotient 3.109 is highlighted in blue in the left hand module, and the submodule 3.111 is highlighted in blue in the right hand module.
We see that there are two copies of $j_{!}\mathcal{O}_{U}$ in this set-up: one as a quotient of the left-hand module $j_{!}f^{s}\mathcal{O}_{U}^{(n)}$, and one as a submodule of a quotient of the right-hand module $j_{+}f^{s}\mathcal{O}_{U}^{(n)}$. These two copies can be naturally identified as follows.
@@
By construction, the map $\overline{\mathrm{can}}$ is injective. Its image is exactly $\ker(s:\Xi_{\rho}\mathcal{O}_{U}\rightarrow\Xi_{\rho}\mathcal{O}_{U})$. This is immediately apparent in Figure 8. Hence $\overline{\mathrm{can}}$ provides an explicit isomorphism which can be used to identify the two copies of $j_{!}\mathcal{O}_{U}$. Under this identification, the algebraic Jantzen filtration 3.110 and the geometric Jantzen filtration 3.112 clearly align.

### v2

@@ 2.4.3. Step 3: !!!-pushforward
As in 2.4.2, this $\widetilde{\mathcal{U}}$-module has an $n$-step filtration by powers of $s$, and decomposes into a direct sum of $\widetilde{\mathcal{U}}$-submodules, each spanned by the set of monomials $\{x_{1}^{k}x_{2}^{\ell}s^{m}\}$ and $\{x_{1}^{a}\partial_{2}^{b}s^{m}\}$ such that $k+\ell=a-b$ is a fixed integer. For $k+\ell=a-b=\lambda$, this submodule is isomorphic to a deformed Verma module of highest weight $\lambda$. We illustrate the module corresponding to $\lambda=0$ in Figure 4
@@ 2.5. The monodromy filtration and the geometric Jantzen filtration
We begin by recalling monodromy filtrations in abelian categories, following [Del80, §1.6]. Given an object $A$ in an abelian category $\mathcal{A}$ and a nilpotent endomorphism $s:A\rightarrow A$, it follows from the Jacobson-Morosov theorem [Del80, Proposition 1.6.1] that there exists a unique increasing exhaustive filtration $\mu^{\bullet}$ on $A$ such that $s\mu^{n}\subset\mu^{n-2}$, and for $k\in\mathbb{N}$, $s^{k}$ induces an isomorphism $\mathrm{gr}_{\mu}^{k}A\simeq\mathrm{gr}_{\mu}^{-k}A.$ This unique filtration is called the monodromy filtration of $A$.
Following Deligne’s proof in [Del80, §1.6], the monodromy filtration can be described explicitly in terms of powers of $s.$ Namely, if we set
@@
to be the image filtration of $A$, then $\mu^{\bullet}$ is the convolution of the kernel and image filtrations; i.e.,
(2.95) $\mu^{r}=\sum_{p-q=r}\mathscr{K}^{p}\cap\mathscr{I}^{q}.$
The monodromy filtration $\mu^{\bullet}$ induces filtrations $J_{!}^{\bullet}$ on $\ker s$ and $J_{+}^{\bullet}$ on $\operatorname{coker}s$. By (2.95), these can be seen to be
@@
(Geometric deformation direction) There are other choices of regular functions on $\widetilde{X}$ which we could have used in the construction of these filtrations. In particular, if $\gamma\in\mathfrak{h}^{*}$ is dominant and integral such that $\gamma(h)=n$ for $n\in\mathbb{Z}_{>0}$, then the function $f_{\gamma}:(x_{1},x_{2})\mapsto x_{2}^{n}$ can be used to define an intermediate extension functor $\Xi_{f_{\gamma}}$ and corresponding Jantzen filtrations. Beilinson–Bernstein establish that all such $f_{\gamma}$ lead to the same filtration. For general Lie algebras $\mathfrak{g}$, the construction can also be done for other choices of meromorphic functions on $\widetilde{X}$, but it is unclear geometrically whether these result in different filtrations [BB93, §4.3]. This is comparable to the dependence on deformation direction in the algebraic Jantzen filtration, see Remark 1.3.
@@ 2.6.1. The algebraic Jantzen filtration
Denote by $I(\lambda)$ the corresponding dual Verma module, defined to be the direct sum of weight spaces in
(2.101) $\operatorname{Hom}_{\mathcal{U}(\overline{\mathfrak{b}})}(\mathcal{U}(\mathfrak{g}),\mathbb{C}_{\lambda}).$
@@
Let $V$ be a $(\mathfrak{g},A)$-bimodule on which the right and left actions of $\mathbb{C}$ agree. The deformed weight space $V^{\mu}$ of $V$ corresponding to a weight $\mu$ is the subspace
(2.103) $V^{\mu}:=\{v\in V\mid(h-\mu(h))v=v\varphi(h)\text{ for all }h\in\mathfrak{h}\}.$
The direct sum of all deformed weight spaces of $V$ is a $(\mathfrak{g},A)$-submodule of $V$ [Soe08, §2.3].
@@
trivially to $\mathfrak{b}$. Here $h\in\mathfrak{h}$, $a\in A$, and $\varphi$ is as in (2.102). The deformed Verma module $M_{A}(\lambda)$ is equal to the direct sum of its deformed weight spaces.
The deformed dual Verma module $I_{A}(\lambda)$ corresponding to $\lambda$ is the direct sum of deformed weight spaces in the $(\mathfrak{g},A)$-bimodule
(2.106) $\operatorname{Hom}_{\mathcal{U}(\bar{\mathfrak{b}})}(\mathcal{U}(\mathfrak{g}),A_{\lambda}).$
There is a canonical isomorphism [Soe08, Proposition 2.12]
@@
Under this isomorphism, $1\in A$ distinguishes a canonical $(\mathfrak{g},A)$-bimodule homomorphism
(2.108) $\psi_{A,\lambda}:M_{A}(\lambda)\rightarrow I_{A}(\lambda).$
@@
Pulling back the filtration above along the canonical map $\psi_{A,\lambda}$ (2.108) gives a $(\mathfrak{g},A)$-bimodule filtration of $M_{A}(\lambda)$.
@@
$M_{A}(\lambda)^{i}:=\left\{m\in M_{A}(\lambda)\mid\psi_{A,\lambda}(m)\in s^{i}I_{A}(\lambda)\right\},$
where $\psi_{A,\lambda}$ is the canonical map (2.108). By applying the reduction map (2.110) to the filtration layers, we obtain a filtration $M(\lambda)^{\bullet}$ of $M(\lambda)$.
@@ 2.6.2. Relationship between algebraic and geometric Jantzen filtrations
As illustrated in Figures 3 and 4 , the global sections of $j_{!}f^{s}\mathcal{O}_{U}^{(n)}$ and $j_{+}f^{s}\mathcal{O}_{U}^{(n)}$ decompose into direct sums of deformed dual Verma and Verma modules, respectively. [^14] To be more precise, the submodules of $\Gamma(\widetilde{X},j_{!}f^{s}\mathcal{O}_{U}^{(n)})$ and $\Gamma(\widetilde{X},j_{+}f^{s}\mathcal{O}_{U}^{(n)})$ corresponding to an integer $\lambda$ are truncated versions of $M_{A}(\lambda)$ (2.106) and $I_{A}(\lambda)$ (2.106) obtained by taking a quotient so that $s^{n}=0$.. The global sections of $\mathrm{can}$ are the direct sum of $\psi_{A,\lambda}$ (2.108) for all integral $\lambda$. There are two natural filtrations of $j_{!}\mathcal{O}_{U}$ which we have described using this set-up.
@@
This is the geometric Jantzen filtration. It can be realized explicitly in terms of the image of powers of $s$ using 2.96. On global sections, this gives
@@
The map $\mathrm{can}$ is described on basis elements in (2.86). Computing these actions for $\lambda=0$, we see in Figure 8 that $\mathrm{can}$ fixes the right-most column and sends any other monomial on the left to a linear combination of monomials directly above the corresponding monomial on the right. The image of $s_{1}(n)=s\circ\mathrm{can}$ (2.89) is highlighted in grey. The quotient by this image is the maximal extension, which is outlined in the red box. The quotient 2.111 is highlighted in blue in the left hand module, and the submodule 2.113 is highlighted in blue in the right hand module.
We see that there are two copies of $j_{!}\mathcal{O}_{U}$ (each highlighted in blue in Figure 8) in this set-up: one as a quotient of the left-hand module $j_{!}f^{s}\mathcal{O}_{U}^{(n)}$, and one as a submodule of a quotient of the right-hand module $j_{+}f^{s}\mathcal{O}_{U}^{(n)}$. These two copies can be naturally identified as follows.
@@
By construction, the map $\overline{\mathrm{can}}$ is injective. Its image is exactly $\ker(s:\Xi_{\rho}\mathcal{O}_{U}\rightarrow\Xi_{\rho}\mathcal{O}_{U})$. This is immediately apparent in Figure 8. Hence $\overline{\mathrm{can}}$ provides an explicit isomorphism which can be used to identify the two copies of $j_{!}\mathcal{O}_{U}$. Under this identification, the algebraic Jantzen filtration 2.112 and the geometric Jantzen filtration 2.114 clearly align.